ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

8. Analysis

1991 MS Classification: 30,35 Astala, Kari, University of Jyväskylä, Finland Analytic aspects of quasiconformality

A number of different topics in geometric analysis make use of quasiconformal phenomena, i. e. properies of mappings, spaces etc. that behave quantitatively like the corresponding conformal notions but are often much more flexible to use. The areas of applications include diverse fields ranging e. g. from complex dynamics and hyperbolic geometry to partial differential equations and singular integrals. The present talk reviews some of the recent advances in the theory of quasiconformal mappings with emphasis on the analytic aspects of the developments.

ICM 1998 Berlin · Plenary and Invited Lectures